IF 6B+20= A,AND 4B+30=A,THEN=
Answer by ramkikk66(644)

IF 6B+20= A,AND 4B+30=A,THEN=

Then
6B + 20 = 4B + 30 (both are equal to A)
2B = 10
B = 5
A = 6B + 20 = 50
:)